# Generating non-Clifford gate operations through exact mapping between Majorana fermions and $\mathbb{Z}_{4}$ parafermions

### v2

@@ Acknowledgements.
This work was supported by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) through project No. EC221010.
Appendix A Majorana Parity has $1,-1$ Eigenvalues
Let $P$ be a unitary $2\times 2$ matrix that squares non-trivially to identity (the Majorana parity $P_{jk}^{(MF)}$ is an example of such an operator). Since $U$ is unitary, it has a complete set of eigenstates ($2$ eigenstates for a $2\times 2$ matrix). We can explicitly construct the two orthogonal eigenstates of $P$ as follows. Let |$\ket{x}$ be a generic state that is not an eigenstate of $P$. Then, |$\ket{1}$=|$\ket{x}+P\ket{x}$ is an eigenstate of $P$ with an eigenvalue of $1$ and |$\ket{-1}$=|$\ket{x}-P\ket{x}$ is an eigenstate of P with eigenvalue of $-1$. Note that |±$\ket{\pm 1}\neq 0$ since |$\ket{x}$ is not an eigenstate of $P$. Moreover, they are orthogonal to each other. ∎
Appendix B $\mathbb{Z}_{4}$ Parafermion Parity has $1,i,-1,-i$ Eigenvalues
@@
where $c$ and $c’$ ’ are constants. Given that $U$ is unitary, the two equations imply that :
@@
Consider a basis of $4$ dimensional operators consisting of all linearly independent multiplicative combinations of $\psi_{1}$ and $\psi_{2}$.
Let us further assume that $U$ contains an identity term. In this case, Eq. (52) and the fact that $\psi_{1}^{4}=\psi_{2}^{4}=I$ imply that $U$ contains $A=s(I+c\psi_{1}^{3}\psi_{2}+c^{2}\psi_{1}^{2}\psi_{2}^{2}+c^{3}\psi_{1}\psi_{2}^{3})$ and $c^{4}=1$. Assuming $U=A$ and Substituting in Eq. (53) gives
@@ Appendix D Braiding of Majorana fermions under a conserved total parity
In a system of $6$ Majorana fermions where the total parity is conserved to $1$, i.e. the states are restricted to the $+1$ eigenstates of the total parity, the action of the total parity operator is identical to the identity. Hence,
@@
If the system is instead projected on the $-1$ subspace of the total parity, Eq. (56) becomes
@@ Appendix E Braiding of parafermions under a conserved total parity
In a system of $4$ $\mathbb{Z}_{4}$ parafermions where the total parity is conserved to $1$, i.e. the states are restricted to the $+1$ eigenstates of the total parity, the action of the total parity operator is identical to the identity. Hence,

### v3

@@ Acknowledgements.
This work was supported by the Deanship of Research Oversight and Coordination (DROC) at King Fahd University of Petroleum & Minerals (KFUPM) through project No. EC221010, as well as by the Deanship of Research Oversight and Coordination (DROC) and the Interdisciplinary Research Center(IRC) for Intelligent Secure Systems (ISS) at King Fahd University of Petroleum & Minerals (KFUPM) through internal research grant No. INSS2507.
Appendix A Majorana Parity has 1,11,-1 Eigenvalues
[…]
Appendix B $\mathbb{Z}_{4}$ Parafermion Parity has 1,i,1,i1,i,-1,-i Eigenvalues
@@
where $c$ and $c$’ are constants. Given that $U$ is unitary, the two equations imply that :
@@
Consider a basis of $4$ dimensional operators consisting of all linearly independent multiplicative combinations of $\psi_{1}$ and $\psi_{2}$. Let us further assume that $U$ contains an identity term. In this case, Eq. (52) and the fact that $\psi_{1}^{4}=\psi_{2}^{4}=I$ imply that $U$ contains $A=s(I+c\psi_{1}^{3}\psi_{2}+c^{2}\psi_{1}^{2}\psi_{2}^{2}+c^{3}\psi_{1}\psi_{2}^{3})$ and $c^{4}=1$. Assuming $U=A$ and Substituting in Eq. (53) gives
@@ Appendix D Braiding of Majorana fermions under a conserved total parity
In a system of six Majorana fermions where the total parity is conserved to $1$, i.e. the states are restricted to the +1+1 eigenstates of the total parity, the action of the total parity operator is identical to the identity. Hence,
@@
If the system is instead projected on the 1-1 subspace of the total parity, Eq. (56) becomes
@@ Appendix E Braiding of parafermions under a conserved total parity
In a system of four $\mathbb{Z}_{4}$ parafermions where the total parity is conserved to $1$, i.e. the states are restricted to the +1+1 eigenstates of the total parity, the action of the total parity operator is identical to the identity. Hence,